5. (15 points) Calculate Z, HR and SR for carbon dioxide at 325 K and 60 bar by the Lee-Kesler correlations.
Answer #1


Solution- @ Calculation of z whore I- 35k p = Gobar from table of characteristics Propeuties of puse Species- for Carbon di6 from table of values of HEHR I RTC RTC at Py = 0.8126 2 (HR) - Tr=1.068 ERIC - 0.958 at Pr = 0.81262 (HR) To - 1.068 RTE No